i tried this problem, but i am stuck. the problem is as follows:
the length of a rectangle is 5cm more than twice its width. find the largest possible width if the perimeter is at most 64cm.
Thank you so much for your help. i really appreciate it = )

Answer by checkley71(8403)

LENGTH=2W+5
WIDTH=W
2(2W+5)+2W<=64
4W+10+2W<=64
6W<=64-10
6W<=54
W<=54/6
W<=9 ANSWER.
L=2*9+5
L=18=5
L=23 ANSWER.
PROOF
2(2*9+5)+2*9<=64
2(18+5)+18<=64
2*23+18<=64
46+18<=64
64<=64